# IndiBias: A Benchmark Dataset to Measure Social Biases in Language Models for Indian Context

### v1

@@ Abstract
The pervasive influence of social biases in language data has sparked the need for benchmark datasets that capture and evaluate these biases in Large Language Models (LLMs). Existing efforts predominantly focus on English language and the Western context, leaving a void for a reliable dataset that encapsulates India’s unique socio-cultural nuances. To bridge this gap, we introduce IndiBias, a comprehensive benchmarking dataset designed specifically for evaluating social biases in the Indian context. We filter and translate the existing CrowS-Pairs dataset to create a benchmark dataset suited to the Indian context in Hindi language. Additionally, we leverage LLMs including ChatGPT and InstructGPT to augment our dataset with diverse societal biases and stereotypes prevalent in India. The included bias dimensions encompass gender, religion, caste, age, region, physical appearance, and occupation. We also build a resource to address intersectional biases along three intersectional dimensions. Our dataset contains 800 filtered sentences from the CrowS-Pairs dataset and tuples for bias measurement across different demographics. It is made available in English and Hindi languages, providing a size comparable to existing benchmark datasets. Furthermore, using IndiBias we compare ten different language models on multiple bias measurement metrics. We observed that the language models exhibit more bias across a majority of the intersectional groups. All the scripts and datasets utilized in this study will be made publicly available. $\ddagger$ Work done at IIT Bombay. This work has been accepted at NAACL 2024.
@@
Nihar Ranjan Sahoo, Pranamya Prashant Kulkarni, Narjis Asad, Arif Ahmad, Tanu Goyal, Aparna Garimella, Pushpak Bhattacharyya Indian Institute of Technology Bombay, India Google, India, Adobe Research, India {nihar, narjisasad, pb}@cse.iitb.ac.in, pranamyakulkarni@gmail.com arifahmadpeace@gmail.com, tanugoyal@google.com, garimell@adobe.com
@@ 1 Introduction
Language models (LMs) are trained on vast amounts of text data and excel in various natural language processing (NLP) tasks. However, many recent studies have shown evidence of undesirable biases and stereotypes in NLP datasets and models Blodgett et al. (2020); Bender et al. (2021). These models stand a risk of reproducing the learned harmful biases in downstream NLP applications Savoldi et al. (2021); Ziems et al. (2022); Mozafari et al. (2020) which in turn can be significantly detrimental to certain demographic groups. This necessitates the need for high-quality benchmark datasets to measure models’ preference for stereotypical associations in diverse social contexts.
@@
    A dataset of 1600 sentences (800 English and 800 Hindi) obtained by translating the Crows-Pairs dataset and changing culture-specific terms in the translation from the US context to the Indian context (section 4.1).
@@ 2 Related Work
Bias in LMs refers to the presence of unfair or discriminatory behavior exhibited by these models towards certain demographic groups or sensitive topics Hammersley and Gomm (1997). Numerous studies demonstrate that LMs tend to reflect and amplify societal biases present in the pre-training data Bolukbasi et al. (2016); Jia et al. (2020); Zhao et al. (2017); Sheng et al. (2021).
While most efforts to detect and mitigate bias in LMs focus on the English language and Western society, recent works address biases in data and language representations from diverse cultures and languages like Arabic Lauscher et al. (2020), French Kurpicz-Briki (2020), Italian Sanguinetti et al. (2020), Japanese Takeshita et al. (2020), etc. There are also few studies addressing this problem in the Indian context Pujari et al. (2020); Malik et al. (2022).
@@
Furthermore, previous works have primarily focused on biases along a single dimension, such as gender or race, with limited exploration of intersectional biases across multiple relevant characteristics. Tan and Celis (2019) propose methods to examine intersectional identities specifically along the race-gender axis using BERT and GPT-2 models, with a focus on African American and European American racial groups. In contrast, our work aims to evaluate the extent of intersectional biases in LMs trained on a multilingual corpus that includes various Indian languages across three intersectional axes, i.e, gender-religion, gender-caste, gender-age.
@@ 3 Characterization of Social Biases in Indian Context
Axes of disparities like Caste, Religion, and Region exhibit a rich array of social biases specific to the Indian context. Caste-based prejudices have a long-standing prevalence in India, despite the efforts for their eradication by the Indian society Ambedkar (2014). The Indian entertainment media, too, has highlighted the plight of sections of society that are at the receiving end of caste and class-based discrimination with movies like Article 15 (2019), The Kashmir Files (2022) [^1] https://rb.gy/1m12wx, https://rb.gy/ut7ggo, and Masaan (2015). Dalits, Adivasis, Denotified Tribes [^2] https://rb.gy/032glh,https://rb.gy/rt002e, and women in backward regions in India face myriad societal biases and detestable stereotypes. Early work by de Souza (1977) highlighted the presence of various stereotypes for regional subgroups in India, by demonstrating an association of character traits with people’s regional identities. Bhatt et al. (2022) conform with de Souza’s work by demonstrating similar associations in datasets- Wikipedia and IndicCorp-en corpus and LMs- MuRIL and mBERT. Works like Rajadesingan et al. (2019), Haokip (2021), Sabharwal and Sonalkar (2015), and McDuie-Ra (2012) also make significant contributions towards spotlighting the specific biases and stereotypes faced by groups of individuals in the Indian society.
Moreover, social biases and stereotypes have a multi-fold nature, possessing global and geo-cultural context-specific elements. Some global axes of social disparities are Gender, Age, and Physical Appearance. However, these global axes too exhibit a variation across different demographics. For instance, consider Gender, an axis of disparity that sees various commonly experienced biases and stereotypes by the fairer sex. However, there are also geo-cultural context-specific biases against women which may exhibit a vast amount of variation across the globe. To illustrate this more clearly, consider the following example sentences:
@@
Figure 1: Examples of paired instances (sent_more-sent_less) from Indian Crows-pairs corpus. Both the examples mentioned here are of stereo type. For more examples, refer to the table 4 in the Appendix.
@@ 4 IndiBias Dataset
We take a multifaceted approach to create IndiBias. It is a composition of modified sentence pairs from CrowS-Pairs (an existing well-known benchmarking dataset) adapted to the Indian context and template-based sentences created by leveraging the power of LLMs. The following subsections provide a detailed description of the dataset creation process.
4.1 Indian CrowS-Pairs
@@
CrowS-Pairs addresses social biases corresponding to nine categories. Sentence pairs corresponding to race, religion, and nationality categories were not relevant to the Indian context. Also, on manual analysis, we found the sentence pairs corresponding to sexual orientation category barely applicable to the Indian context. Many of the phrases related to sexual orientation from the original dataset do not have a proper translation in Hindi. So, we first filtered CrowS-Pairs to retain sentences corresponding to gender, age, disability, physical appearance, and socioeconomic status categories. We retain those sentence pairs that are pertinent to the Indian context or can be slightly adapted to align with the Indian context. This filtered subset contains 800 sentence pairs, out of which the bias categories of gender, age, disability, physical appearance, and socioeconomic status had a share of $43.5\%,13.5\%,6.7\%,9.9\%,$ and $26.4\%$ respectively.
The second step was to obtain preliminary machine translations of sentences in this subset; we used NLLB Translator [^3] https://rb.gy/zo71s5 and Google Translate [^4] https://pypi.org/project/googletrans/ for this. Five annotators meticulously reviewed the original sentence pairs and their Hindi translations obtained in the previous step. The translation best suited was selected. In many instances, manual modifications were also done to rectify any machine translation errors. To eliminate annotator bias while selecting a translation, information regarding the machine translation engine, i.e., Google Translate and NLLB, was kept hidden from them. Sentence pairs containing named entities were replaced with the most common Indian male and female names obtained from resources discussed in section 5.1. Each sentence pair in the Indian version of CrowS-Pairs contains two sentences, S1 and S2. Every sentence contains a target group and an attribute that may be stereotypically associated with the target group. The two sentences in a pair satisfy the property of being minimally distant. Two sentences are said to be minimally distant if the only words they differ in are the target group and related pronouns. For example: S1: He was excited about his big football game tomorrow. S2: She was excited about her big football game tomorrow.
In our dataset, pairs in which S1 represents a commonly held stereotype and S2 is a less stereotyping sentence are labeled stereo. The pairs in which S1 is a less stereotyping sentence and S2 is a commonly held stereotype are labeled anti-stereo. Challenges in adapting the existing sentence pairs to the Indian context are discussed in 4.3.
@@
4.2 Bias tuple creation
The axes of region and caste being specific to the Indian context are absent in the original CrowS-Pairs dataset. To address this, we augment our dataset by creating tuples designed specifically for the Indian context. The tuples created encompass axes of Region, Caste, Religion, Age, Gender, Physical appearance, and occupation/socioeconomic status. This also makes the dataset more exhaustive by capturing the prevalent stereotypes and biases that may be lacking in the Crows-Pairs. This tuple dataset can be used to further extend the Indian CrowS-Pairs dataset. The tuples follow the following format: (identity term, stereotypical attribute). Where identity term represents a specific group, and attribute is a concept stereotypically associated with the identity term. The identity terms included are listed in figure 2. A tuple is characterized as a positive tuple if the attribute describing the identity term has a positive connotation and is otherwise characterized as a negative tuple. For tuple creation, we follow a four-step process. We first prompt ChatGPT/InstructGPT to generate ten positive and ten negative attributes for each of the included identity terms. The specific prompts used can be found in Table 5 of the Appendix C. Next, three annotators are employed to evaluate whether the identity term and attribute tuples reflect prevalent stereotypical associations in Indian society. Tuples marked as stereotypical by $\geq$ 2 annotators are added to the dataset. Examples of the selected tuples, along with the number of annotators who labeled them as stereotypical, and the corresponding type i.e., positive or negative, are provided in table 7 in the Appendix F. Our approach to tuple generation differs from Bhatt et al. (2022), which captured only 2 bias axes, namely region and religion. We capture 5 additional axes and make the tuples available in the Hindi language also.
Details of the annotation procedure and the inter-annotator scores are discussed in Appendix B.
@@ 4.3 Challenges in Dataset Creation
Machine Translation and Target Language Properties: There were numerous instances where the machine translations were either incorrect or did not appropriately represent the desired intent of the source sentence. The annotators modified such translations suitably. There were many sentence pairs with identical Hindi translations for both S1 and S2. This phenomenon was observed because, in Hindi, pronouns are not gendered. Hence source sentences that differed only in words ‘he’ and ‘she’, ‘his’ and ‘her’ were found to be identical post-translation. Words he and she both being translated to ‘vah’, and his and hers both translated to ‘uska/unka.’ To retain gender information post-translation we modified the sentences to include phrases like ’ek purush (a man)/ ek mahila (a woman)’. Figure 1(b) contains an example to demonstrate this challenge. More examples that demonstrate some innovative resolutions we provided for this problem are included in figure 4 in the Appendix E.
@@
Adapting sentences to the Indian Context: Sentences containing phrases like ’rural Kentucky’, ’star-quarterback’, etc have little to no relevance to the Indian context. These were modified suitably, see example Figure 1(a). Also, assuring that the Hindi translations of source sentences are befitting to reflect commonly held stereotypes by Indian society was another major challenge.
@@
We also consider the pitfalls discussed by Blodgett et al. (2021) and address many of them while creating the dataset. We discuss our approach to addressing those pitfalls in table 6 of Appendix E.

### v2

@@ Abstract
The pervasive influence of social biases in language data has sparked the need for benchmark datasets that capture and evaluate these biases in Large Language Models (LLMs). Existing efforts predominantly focus on English language and the Western context, leaving a void for a reliable dataset that encapsulates India’s unique socio-cultural nuances. To bridge this gap, we introduce IndiBias, a comprehensive benchmarking dataset designed specifically for evaluating social biases in the Indian context. We filter and translate the existing CrowS-Pairs dataset to create a benchmark dataset suited to the Indian context in Hindi language. Additionally, we leverage LLMs including ChatGPT and InstructGPT to augment our dataset with diverse societal biases and stereotypes prevalent in India. The included bias dimensions encompass gender, religion, caste, age, region, physical appearance, and occupation. We also build a resource to address intersectional biases along three intersectional dimensions. Our dataset contains 800 sentence pairs and 300 tuples for bias measurement across different demographics. The dataset is available in English and Hindi, providing a size comparable to existing benchmark datasets. Furthermore, using IndiBias we compare ten different language models on multiple bias measurement metrics. We observed that the language models exhibit more bias across a majority of the intersectional groups. All the scripts utilized and datasets created in this study are publicly available [^1] https://github.com/sahoonihar/IndiBias. $\ddagger$ Work done at IIT Bombay. This work has been accepted at NAACL 2024.
@@
Nihar Ranjan Sahoo, Pranamya Prashant Kulkarni, Narjis Asad, Arif Ahmad, Tanu Goyal, Aparna Garimella, Pushpak Bhattacharyya IIT Bombay, India, Google, India, Adobe Research, India {nihar, narjisasad, pb}@cse.iitb.ac.in, pranamyakulkarni@gmail.com arifahmadpeace@gmail.com, tanugoyal@google.com, garimell@adobe.com
@@ 1 Introduction
Language models (LMs) are trained on vast amounts of text data and excel in various natural language processing (NLP) tasks. However, many recent studies have shown evidence of undesirable biases and stereotypes in NLP datasets and models Blodgett et al. (2020); Bender et al. (2021); Sahoo et al. (2022). These models stand a risk of reproducing the learned harmful biases in various downstream NLP applications Savoldi et al. (2021); Ziems et al. (2022); Mozafari et al. (2020) which in turn can be significantly detrimental to certain demographic groups. This necessitates the need for high-quality benchmark datasets to measure models’ preference for stereotypical associations in diverse social contexts.
@@
    A dataset of 1600 sentence pairs (800 English and 800 Hindi) obtained by translating the Crows-Pairs dataset and changing culture-specific terms in the translation from the US context to the Indian context and also by leveraging tuple dataset (section 4.2).
@@ 2 Related Work
Bias in LMs refers to the presence of unfair or discriminatory behavior exhibited by these models towards certain demographic groups or sensitive topics Hammersley and Gomm (1997); Singh et al. (2022). Numerous studies demonstrate that LMs tend to reflect and amplify societal biases present in the pre-training data Bolukbasi et al. (2016); Jia et al. (2020); Zhao et al. (2017); Sheng et al. (2021).
While most efforts to detect and mitigate bias in LMs focus on the English language and Western society, recent works address biases in data and language representations from diverse cultures and languages like Arabic Lauscher et al. (2020), French Kurpicz-Briki (2020), Italian Sanguinetti et al. (2020), etc. There are also few studies addressing this problem in the Indian context Pujari et al. (2020); Malik et al. (2022).
@@
Furthermore, previous works have primarily focused on biases along a single dimension with limited exploration of intersectional biases across multiple relevant characteristics. Tan and Celis (2019) propose methods to examine intersectional identities specifically along the race-gender axis using BERT and GPT-2 models, with a focus on African American and European American racial groups. In contrast, our work aims to evaluate the extent of intersectional biases in LMs trained on a multilingual corpus that includes various Indian languages across three intersectional axes, i.e, gender-religion, gender-caste, gender-age.
@@
Figure 1: Examples of paired instances (S1-S2) from Indian Crows-pairs (ICS) corpus. Both the examples mentioned here are of stereo type. S1 always presents a stereotype or an anti-stereotype for the corresponding bias type. The Hindi examples mentioned here are the Hindi versions of the corresponding Modified (English) pair. Construction of sentence pairs and issues mentioned in the Concern column are elaborated in sections 4.2.1, 4.3. For more examples, refer to the table 6, 7 in the Appendix.
@@ 3 Characterization of Social Biases in Indian Context
Axes of disparities like Caste, Religion, and Region exhibit a rich array of social biases specific to the Indian context. Caste-based prejudices have a long-standing prevalence in India, despite the efforts for their eradication by the Indian society Ambedkar (2014). The Indian entertainment media, too, has highlighted the plight of sections of society that are at the receiving end of caste and class-based discrimination with movies like Article 15 (2019), The Kashmir Files (2022) [^2] https://rb.gy/1m12wx, https://rb.gy/ut7ggo, and Masaan (2015). Dalits, Adivasis, Denotified Tribes [^3] https://rb.gy/032glh,https://rb.gy/rt002e, and women in backward regions in India face myriad societal biases and detestable stereotypes. Early work by de Souza (1977) highlighted the presence of various stereotypes for regional subgroups in India, by demonstrating an association of character traits with people’s regional identities. Bhatt et al. (2022) conform with de Souza’s work by demonstrating similar associations in datasets- Wikipedia and IndicCorp-en corpus and LMs- MuRIL and mBERT. Works like Sahoo et al. (2023), Rajadesingan et al. (2019), Haokip (2021), Sabharwal and Sonalkar (2015), and McDuie-Ra (2012) also make significant contributions towards spotlighting the specific biases and stereotypes faced by groups of individuals in the Indian society.
Moreover, social biases and stereotypes have a multi-fold nature, possessing global and geo-cultural context-specific elements. Some global axes of social disparities are Gender, Age, and Physical Appearance. However, these global axes too exhibit a variation across different demographics. For instance, consider Gender, an axis of disparity that sees various commonly experienced biases and stereotypes by women. However, there are also geo-cultural context-specific biases against women which may exhibit a vast amount of variation across the globe. To illustrate this more clearly, consider the following example sentences:
@@ 4 IndiBias Dataset
We take a multifaceted approach to create IndiBias. It is a composition of modified sentence pairs from CrowS-Pairs (an existing benchmarking dataset) adapted to the Indian context, sentences generated using IndiBias tuples, and template-based sentences created by leveraging the power of LLMs. The following subsections provide a detailed description of the dataset creation process.
4.1 IndiBias: Bias tuples
The axes of region and caste being specific to the Indian context are absent in the original CrowS-Pairs dataset. To address this, we created tuples designed specifically for the Indian context. The tuples created encompass axes of Region, Caste, Religion, Age, Gender, Physical appearance, and occupation/socioeconomic status. This also makes the dataset more exhaustive by capturing the prevalent stereotypes and biases that may be lacking in the Crows-Pairs. We also use these tuples to further extend the India Crow-Pairs dataset as described in section 4.2.2. The tuples are in the following format: (identity term, stereotypical attribute). Where identity term represents a specific group, and attribute is a concept stereotypically associated with the identity term. The identity terms included are listed in figure 2. A tuple is characterized as a positive tuple if the attribute describing the identity term has a positive connotation and is otherwise characterized as a negative tuple. For tuple creation, we follow a four-step process. We first prompt ChatGPT/InstructGPT to generate 10 positive and 10 negative attributes for each of the included identity terms. The specific prompts used can be found in Table 5 of the Appendix F. Next, three annotators are employed to evaluate whether the identity term and attribute tuples reflect prevalent stereotypical associations in Indian society. Tuples marked as stereotypical by $\geq$ 2 annotators are considered stereotypical pairs. Examples of the selected tuples, along with the number of annotators who labeled them as stereotypical, and the corresponding type, i.e., positive or negative, are provided in table 8 (Appendix I). Our approach to tuple generation differs from Bhatt et al. (2022), which captured only 2 bias axes, namely region and religion. We capture 5 additional axes and also use a human-LLM partnership approach to generate stereotypical sentence pairs using these tuples. Details of the annotation procedure and the inter-annotator scores are discussed in Appendix D.
4.2 IndiBias: Indian CrowS-Pairs (ICS)
We created a Crow-Pairs style dataset explicitly tailored to the Indian socio-cultural landscape to assess biases in multilingual LLMs across seven distinct social bias axes: gender, religion, age, caste, disability, physical appearance, and socioeconomic status. Initially, we filtered and adapted the original Crows-Pairs to align with the nuances of the Indian context. Subsequently, we augmented that with a similar dataset, exploiting the IndiBias tuple dataset.
4.2.1 Inherited from Nangia et al. (2020)
@@ 4.2 IndiBias: Indian CrowS-Pairs (ICS)
CrowS-Pairs addresses social biases corresponding to nine categories. Sentence pairs corresponding to race [^4] Race category pairs were mostly for Black vs White, which is not applicable to Indian society., religion [^5] Pairs mostly centered on Christian and Jewish communities, which aren’t prevalent biases in India. Pairs containing Muslim groups also lack reflection of Indian societal biases., and nationality [^6] Pairs in it compare attributes among countries; however, for our dataset, we only focus on Indian society. categories were not relevant to the Indian context. Also, on manual analysis, we found the sentence pairs corresponding to sexual orientation category barely applicable to the Indian context. Many of the phrases related to sexual orientation from the original dataset do not have a proper translation in Hindi. So, we first filtered CrowS-Pairs to retain sentences corresponding to gender, age, disability, physical appearance, and socioeconomic status categories. We retain those sentence pairs that are pertinent to the Indian context or can be modified and adapted to align with the Indian context. This filtered subset contains 542 sentence pairs, out of which the bias categories of gender, age, disability, physical appearance, and socioeconomic status has a share of $45.6\%,14.4\%,5.5\%,9.6\%,$ and $24.9\%$ respectively; for more details refer Table 6.
@@
The second step was to obtain preliminary machine translations of sentences in this subset; we used NLLB Translator [^7] https://rb.gy/zo71s5 and Google Translate [^8] https://pypi.org/project/googletrans/ for this. Five annotators meticulously reviewed the original sentence pairs and their Hindi translations obtained in the previous step. The translation best suited was selected. In many instances, manual modifications were also done to rectify any machine translation errors. To alleviate annotator bias while selecting a translation, information regarding the machine translation engine, i.e., Google Translate and NLLB, was kept hidden from them. Sentence pairs containing named entities were replaced with the most common Indian male and female names obtained from resources discussed in section 5.1. Each sentence pair in the Indian version of CrowS-Pairs contains two sentences, S1 and S2. Every sentence contains a target group and an attribute that may be stereotypically associated with the target group. The two sentences in a pair satisfy the property of being minimally distant. Two sentences are said to be minimally distant if the only words they differ in are the target group and related pronouns. For example: S1: He was excited about his big football game tomorrow. S2: She was excited about her big football game tomorrow. We remove or modify the pairs from Crows-Pairs when both attribute terms and target groups differ between two participating sentences.
4.2.2 Using IndiBias Tuples
We employ a human-LLM partnership approach to create similar sentence pairs as described in 4.2.1. Section 4.1 describes the method used for obtaining tuples. The tuples that were marked as stereotypical by $\geq$ 2 annotators were used to generate stereotypical sentences by prompting LLMs namely ChatGPT and Gemini. Given a tuple of the form (identity, attribute), LLM was prompted to generate naturally occurring sentences that demonstrate the identity group being associated stereotypically with the attribute in the tuple. A naturally occurring sentence is one that occurs in social media platforms, novels, human conversations, movie dialogues, etc. The prompts used for generating these sentences are listed in table 5. The LLM-generated sentences were verified, filtered, and modified by 4 human annotators. Suitable sentences for which meaningful sentence pairs could be created were selected. The sentence pairs obtained using dataset by Nangia et al. (2020) address categories gender, age, disability, physical appearance, and socioeconomic status, while creating sentences from tuples, we focused on two categories, namely religion and caste. The created sentence pairs were then translated to parallel Hindi pairs using the same approach discussed in section 4.2.1. Some examples of the stereotypical tuples and corresponding sentence pairs are given in table 7.
We created a total of 258 sentence pairs using tuples, out of which the bias categories of religion and caste have a share of 62.6%, 37.4% respectively999for more details regarding dataset refer Table 6. Our work on the creation of IndiBias Tuples aligns with Jha et al. (2023) and Bhatt et al. (2022). However, to the best of our knowledge, no other work so far has employed this unique human-LLM partnership approach to generate sentence pairs for assessing the presence of learned stereotypes in language models using stereotypical tuples. We also made sure to meticulously avoid all the pitfalls outlined in table 7.
In our dataset, a sentence pair is labeled stereo when the target group in S1 has a stereotypical association with the attribute in S1. It is labeled as antistereo when the attribute in S1 negates or represents the opposite of the actual stereotype (anti-stereotype) associated with the target group in S1. Challenges in adapting the existing sentence pairs to the Indian context are discussed below.
@@ 4.3 Challenges in Dataset Creation
We divide the challenges in adapting the existing sentences from CrowS-Pairs to the Indian context in four broad categories. We also consider the pitfalls discussed by Blodgett et al. (2021) and address many of them while creating the dataset. We discuss our approach to addressing those pitfalls in table 7 of Appendix H.
Machine Translation and Target Language Properties: There were numerous instances where the machine translations were either incorrect or did not appropriately represent the desired intent of the source sentence. The annotators modified such translations suitably. There were many sentence pairs with identical Hindi translations for both S1 and S2. This phenomenon was observed because, in Hindi, pronouns are not gendered. Hence source sentences that differed only in words ‘he’ and ‘she’, ‘his’ and ‘her’ were found to be identical post-translation. Words he and she both being translated to ‘vah’, and his and hers both translated to ‘uska/unka.’ To retain gender information post-translation we modified the sentences to include phrases like ‘ek purush (a man)/ ek mahila (a woman)’. Figure 1(b) contains an example to demonstrate this challenge. More examples that demonstrate some innovative resolutions we provided for this problem are included in figure 6 in the Appendix H.
@@
Adapting sentences to the Indian Context: Sentences containing phrases like ‘rural Kentucky’, ‘star-quarterback’, etc have little to no relevance to the Indian context. These were modified suitably, see example Figure 1(a). Also, assuring that the Hindi translations of source sentences are befitting to reflect commonly held stereotypes by Indian society was another major challenge.